Two charges +a and +4q are separated by a distance 2r. Use Gauss' law to find the electric field at a distance. R around charge +q
1
Expert's answer
2021-08-24T05:19:02-0400

Due to +q charge flux

"\phi_1=\frac{q}{\epsilon_0}"

Due to +4q charge flux

"\phi_1=\frac{4q}{\epsilon_0}"


Net flux

"\phi=\phi_1+\phi_2"

"\phi=\frac{q}{\epsilon_0}+\frac{4q}{\epsilon_0}=\frac{5q}{\epsilon_0}"

Electric field

"E_1=\frac{kQ}{R^2}\\E_2=\frac{4kq}{R^2}"

"E=E_2-E_1"

"E=\frac{4kq}{R^2}-\frac{kq}{R^2}=\frac{3kq}{R^2}"

Net direction (E.F)of electric field E2 direction